Widget Co. was founded at the start of 2016 by issuing 10M shares at $187.50 per share and $1,200M of long term debt. All data in the table, except actual price per share, is in millions. No assets were sold during the first two years of operation. Remember when and where to include marketable securities and the current portion of long term debt in your intermediate calculations. Use this information and the table below to answer problems 28-30. Net Income Dividends Cash Marketable Sec. Acct. Rec. Inventory Gross PPE cumulated Dep. Net PPE Acct. Payable CP LT Debt LT Debt 2016 250 125 100 100 300 300 3000 200 2800 400 0 1200 100 2017 300 150 120 120 300 360 3600 500 3100 400 100 1400 98 Shares Out $20.0022.00

28. Between December 2016 and December 2017, Widget Company’s market capitalization:

  1. Increased by $332M

  2. Increased by $200M

  3. Increased by $156M

  4. Decreased by $156M

  5. Decreased by $200M

29. At the end of 2016, Widget Company’s P/E is closest to:

a. 7.2 b. 8.0 c. 14.4 d. 15.0 e. 16.0

30. During 2017, Widget Company __________.
HINT: BOY SE + NI – DIVS + Equity Issued – Equity Repurchased = EOY SE 23

  1. Net borrowing was $200M and shares outstanding were repurchased at a price of $24 per share

  2. Net borrowing was $200M and the company issued shares at a price of $25 per share

  3. Net borrowing was $300M and shares outstanding were repurchased at a price of $24 per share

  4. Net borrowing was $300M and shares outstanding were repurchased at a price of $25 per share

  5. Net borrowing was $300M and the company issued shares at a price of $24 per share
Answer #1

1. Market Capitalization= Price/Share * Number of Shares Outstanding

M.Cap(2016)= 100*20=2000

M.Cap(2017)= 98*22=2156

Change in M.Cap=2156-2000=156

2. At the end of 2016, Earnings per share= Net Income/ Shares Outstanding = 250/100=2.5

Price to earnings= Price per share/Earnings per share= 20/2.5= 8.0

3. Net borrowing = LT Debt + CP LT Debt= 100+200=300

As given in Hint, the formula to calculate Equity Issued/Repurchased,

SE(Shareholders Equity)= Assets- Liabilities= Cash+Marketable Securities+Acc. Rec+ Inventory+Net PPE- Acc. Pay. - Debt

BOY SE for 2017= 100+100+300+300+2800-400-1200= 2000

EOY SE for 2017= 120+120+300+360+3100-400-100-1400= 2100

2000+300-150+x=2100

x= -50million

So Equity was repurchased

Change in Equity outstanding= 2 million

$50million/2million= $25 per share